# A user-friendly package and workflow for generating effective homogeneous rheologies for the study of the long-term orbital evolution of multilayered planetary bodies

### 29 Mar 2026 v2

@@ Abstract
Aims. We present a user-friendly, open-source Wolfram Language package that automates the construction of an effective homogeneous generalized Voigt rheology for a spherically symmetric, incompressible layered body with Maxwell solid layers. It provides a practical bridge between layered interior models and time-domain simulations of tidal evolution.
@@
Results. The package returns the parameters of the equivalent homogeneous model, including elastic, gravitational, viscous, and Voigt-element contributions, in a format suitable for downstream numerical applications. As a case study, we apply the package to a five-layer lunar interior model and obtain its equivalent generalized Voigt representation, together with a reduced model that preserves the tidal response over the frequency interval relevant for orbital evolution while using fewer relaxation elements.
Conclusions. This package makes the reduction from stratified viscoelastic interiors to effective homogeneous rheologies reproducible and accessible. It allows physical tidal dissipation models to be used in long-term orbital and spin-evolution studies without having to repeatedly solve the full layered boundary-value problem.
@@ 1 Introduction
Long-term dynamical evolution codes often rely on simplified homogeneous rheologies or empirical prescriptions, such as constant time lag, constant phase lag, or single-element Maxwell models. These approximations are computationally convenient, but generally fail to reproduce the multiple relaxation timescales and dissipation peaks that arise in layered bodies Gevorgyan (2021); Bolmont et al. (2020). In Gevorgyan et al. (2023), the authors showed that, under the assumptions of a simple multilayer rheology, the tidal response of a stratified body can be modeled by a homogeneous body with a sufficiently complex generalized rheology, like a generalized Voigt model. This prompted us to develop a relatively simple workflow that allows one to replace the stratified problem with an effective homogeneous model without losing the frequency-dependent tidal response.
In this work we present an open-source Wolfram Language package [^1] https://github.com/ygevorgyan/HomogeneousModel that takes a user-specified multi-layered model and constructs the corresponding effective homogeneous generalized Voigt rheology. The input file contains the layer radii, densities, shear moduli, viscosities, and forcing frequency, and the main routine performs the full reduction pipeline: computation of $k_{2}(s)$ for the layered model, identification of secular poles and residues, inversion to the equivalent compliance function, and extraction of the generalized Voigt parameters. An optional reduction step identifies dominant relaxation modes over a user-defined frequency interval, yielding a compact model when an exact full representation is unnecessary.
Our implementation was partly motivated by the need to connect detailed geophysical interior models with modern time-domain orbital integrators such as the RheoVolution (de Oliveira et al., 2025a, b), which can evolve deformable bodies with complex homogeneous rheologies but require as input the parameters of an effective rheological model rather than a full multilayer interior description. Our package fills this gap by translating a stratified viscoelastic body into a homogeneous generalized Voigt one directly usable in such simulations.
@@ 2.1 Layered viscoelastic model
Consider a spherically symmetric body of NN incompressible, homogeneous layers, each with outer radius $r_{k}$, density $\rho_{k}$, rigidity $\mu_{k}$, and viscosity $\eta_{k}$. Solid layers are assumed to have Maxwell rheology, and their complex shear modulus in the Laplace domain is
@@
The tidal response is described by the degree-2 Love number $k_{2}(s)$, obtained from the propagator-matrix solution of the governing equations for a stratified body. All layers are treated within a unified propagator formulation, and liquid layers are handled in the small-shear limit, see Subsection $3.$
For a finite multilayer Maxwell body, $k_{2}(s)$ is a rational function of the Laplace variable and admits the mode decomposition
$k_{2}(s)=k_{E}+\sum_{i=1}^{n}\frac{r_{i}}{s-s_{i}},$ (2)
where $k_{E}$ is the elastic Love number, $s_{i}<0$ are the secular poles, $r_{i}$ are the corresponding residues, and $\tau_{i}=-s_{i}^{-1}$ are the relaxation times. Thus, the layered tidal response is fully characterized by the set of parameters $\{k_{E},s_{i},r_{i}\}$.
@@ 2.2 The homogeneous model
The equivalent homogeneous representation of stratified bodies is given by a generalized Voigt rheology in Figure 1. The Love number of the body is (Gevorgyan et al., 2023)
$k_{2}(s)=\frac{3\rm{I}_{\circ}G}{R^{5}}\,C(s),$ (3)
where $\rm{I}_{\circ}$ is the mean moment of inertia, $C(s)$ is the complex compliance, and $R$ is the mean radius of the body. The gravitational rigidity $\gamma$ is defined from the fluid Love number $k_{f}$ as (Correia et al., 2018)
$\gamma=\frac{3\rm{I}_{\circ}G}{R^{5}}\frac{1}{k_{f}}.$ (4)
For the generalized Voigt rheology, the complex compliance is decomposed as
@@
where $\alpha$ is the elastic rigidity and $C_{v}(s)$ contains the viscous contribution of the dashpot and a finite number of Kelvin–Voigt elements in Figure 1. The equivalent homogeneous model is specified by the parameters $\{\gamma,\alpha,\eta,\alpha_{i},\eta_{i}\}$.
2.3 Connection with the software workflow
The software automates the passage from the layered model to the homogeneous one in three steps. First, it computes the layered Love number $k_{2}(s)$ (the tidal Love number computed for the layered interior model) from the propagator formulation. Second, it identifies the secular poles $s_{i}$ and residues $r_{i}$ in the decomposition (2). Third, it constructs the compliance $C(s)$ from equations (3)–(5) and extracts the parameters of the equivalent generalized Voigt model.
The explicit use of the pole–residue representation of the layered Love number as the intermediate object in the inversion pipeline is particularly effective from both a numerical and conceptual standpoint. It provides a compact spectral description of the viscoelastic response that is well suited to numerical computation and allows the equivalent generalized Voigt rheology to be constructed in a physically interpretable way.
@@ 3.2 Input configuration
Table 1: Main user inputs and numerical controls of the Wolfram package. Physical quantities are supplied through ModelInput.wl; optional numerical settings control pole search and dominant-mode selection.
@@
rhok kg m-3 Layers density
@@ 3.3 Main routine and computational workflow
performs the parameter extraction through the sequence of operations illustrated in Fig. 2. It follows the same three-step structure of Sections 1 and 2: forward computation of the layered Love number, mode decomposition, and rheological inversion. The code first evaluates the elastic Love number $k_{E}$ and the complex tidal response $k_{2}(i\omega)$ for the layered body. It then searches for the zeros of the secular determinant on the negative real axis, refines them numerically, and computes the corresponding residues. From these quantities it evaluates the fluid Love number $k_{f}$, the mapping constants $\Theta$, $\gamma$, and $\alpha$, and finally constructs the compliance
$C(s)=\frac{k_{2}(s)}{\Theta-\gamma k_{2}(s)},$ (6)
from which the parameters $\{\eta,\alpha_{i},\eta_{i}\}$ of the Voigt model are extracted.
@@
The forward layered problem is evaluated through a unified $6\times 6$ propagator matrix approach for all layers. Liquid layers are treated by small-shear regularization, such that solid and liquid regions are handled within the same numerical framework. This is an important practical point: the current package does not switch to a separate Clairaut propagator for liquids, but instead uses a single propagation formalism for arbitrary combinations of solid and liquid layers. A true liquid has $\mu=0$, but the $Y$-matrix (Sabadini et al., 2016, Eq. 2.42) contains terms of the form $\mu/r$ in rows 3, 4, and 6. Setting $\mu=0$ exactly would therefore make these rows degenerate, causing the linear solver to fail when computing the propagator
@@
Instead, we assign a very small effective rigidity $\hat{\mu}\approx 10^{-20}\,\mathrm{Pa}$ to liquid layers. This keeps the matrix full-rank and invertible, while making the shear-stress rows effectively zero to machine precision. This value is negligible compared with physical shear moduli, which are typically of order $10^{9}$$10^{11}\,\mathrm{Pa}$, but remains large enough to preserve numerical stability at all reasonable (actually, even at quite stringent) working precisions. This is a standard regularization in viscoelastic normal-mode calculations to avoid introducing a separate $2\times 2$ or $4\times 4$ propagator for liquid layers while yielding numerically identical results.
After the full generalized Voigt model is obtained, the package performs an optional dominant-mode reduction. For each Voigt element, it evaluates the maximum fractional contribution to $\operatorname{Im}[C(i\omega)]$ over a user-defined frequency interval $\omega_{\min},\omega_{\max}]$. Elements whose contribution exceeds the prescribed threshold are kept in the reduced model. This produces a compact, effective rheology targeted to the frequency interval most relevant for the intended dynamical application.
@@ 3.4 Output
The output falls into three groups: the Love numbers of the layered body, the constants used in the layered-to-homogeneous mapping, and the relaxation modes and rheological parameters of the equivalent generalized Voigt model. This organization separates diagnostic quantities from the final effective model. The Love-number and mode outputs are used to check the inversion, and the generalized Voigt parameters are used in downstream applications. The package also generates summary tables, dominant-mode diagnostics, and comparison plots. The outputs can be used for both analysis and reuse. The user can inspect the mode decomposition, compare the full and reduced models, or export the reduced generalized Voigt parameters to an external integrator such as RheoVolution (de Oliveira et al., 2025a, b).
Table 2: Quantities returned by computeHomogeneous. The routine returns both the intermediate mode information of the layered model and the parameters of the equivalent homogeneous generalized Voigt rheology.
@@
prefactor s-2 Mapping prefactor $\Theta=3\rm{I}_{\circ}G/R^{5}$
@@
nPoles integer Number of secular poles $M$
@@
nVoigt integer Number of Kelvin–Voigt elements $N_{V}$
@@ 4 Case study: A five-layer model of the Moon
To test the package, we use the five-layer lunar model of Gevorgyan et al. (2023) to illustrate the equivalence between a multilayer Maxwell body and an effective homogeneous rheology. The model parameters are summarized in Table 3, and its cross-section is illustrated in Fig. 3; the figure was automatically produced by VisualizeLayers.wl module. The Moon is the default example configuration in the package. The package was also tested for Io, Europa, and Enceladus using models and data from Hussmann and Spohn (2004), Wahr et al. (2006), Bierson and Nimmo (2016), and Thomas et al. (2016), although we are only going to report the results for the Moon. The ModelInput file for these additional test cases can be found in the software package repository.
Table 3: Physical parameters for the 5-layer viscoelastic model of the lunar interior taken from Matsuyama et al. (2016, Table 2) and Gevorgyan et al. (2023).

### 03 May 2026 v3

@@ Abstract
Aims. We present a user-friendly, open-source Wolfram Language package that constructs an effective homogeneous generalized Voigt rheology for a spherically symmetric, incompressible layered body with Maxwell solid layers. The package converts the tidal response of a layered interior model into the parameters required by time-domain simulations of tidal evolution.
@@
Results. The package returns the parameters of the equivalent homogeneous model, including elastic, gravitational, viscous, and Voigt-element contributions, in a format that can be used by external tidal-evolution codes. As a case study, we apply the package to a five-layer lunar interior model and obtain its equivalent generalized Voigt representation, together with a reduced model that preserves the tidal response over the frequency interval relevant for orbital evolution while using fewer relaxation elements.
Conclusions. The package provides a reproducible implementation of the reduction from stratified viscoelastic interiors to effective homogeneous rheologies. It allows tidal dissipation from layered models to be included in long-term orbital and spin-evolution calculations without solving the full layered boundary-value problem at each step.
@@ 1 Introduction
Long-term dynamical evolution codes often rely on simplified homogeneous rheologies or empirical prescriptions, such as constant time lag, constant phase lag, or single-element Maxwell models. These approximations are computationally convenient, but generally fail to reproduce the multiple relaxation timescales and dissipation peaks that arise in layered bodies Gevorgyan (2021); Bolmont et al. (2020). In Gevorgyan et al. (2023), the authors showed that, under the assumptions of a simple multilayer rheology, the tidal response of a stratified body can be modeled by a homogeneous body with a sufficiently complex generalized rheology, like a generalized Voigt model. This motivated us to develop a relatively simple workflow that allows one to replace the stratified problem by an effective homogeneous model that preserves the frequency-dependent tidal response.
In this work we present an open-source Wolfram Language package [^1] https://github.com/ygevorgyan/HomogeneousModel that takes a user-specified multilayered model and constructs the corresponding effective homogeneous generalized Voigt rheology. The input file contains the layer radii, densities, shear moduli, viscosities, and forcing frequency, and the main routine performs the full reduction pipeline: computation of $k_{2}(s)$ for the layered model, identification of secular poles and residues, inversion to the equivalent compliance function, and extraction of the generalized Voigt parameters. An optional reduction step identifies dominant relaxation modes over a user-defined frequency interval, yielding a compact model when an exact full representation is unnecessary.
Our implementation was partly motivated by the need to connect detailed geophysical interior models with modern time-domain orbital integrators such as the RheoVolution (de Oliveira et al., 2025a, b), which can evolve deformable bodies with complex homogeneous rheologies but require as input the parameters of an effective rheological model rather than a full multilayer interior description. The package converts the tidal response of a stratified viscoelastic body into the parameters of a homogeneous generalized Voigt model that can be used in such simulations.
@@ 2.1 Layered viscoelastic model
Consider a spherically symmetric body of $N$ incompressible, homogeneous layers, each with outer radius $r_{k}$, density $\rho_{k}$, rigidity $\mu_{k}$, and viscosity $\eta_{k}$. Solid layers are assumed to have Maxwell rheology, and their complex shear modulus in the Laplace domain is
@@
The tidal response is described by the degree-2 Love number $k_{2}(s)$, obtained from the propagator-matrix solution of the governing equations for a stratified body. All layers are treated within a unified propagator formulation, and liquid layers are handled in the small-shear limit, see Subsection 3.3.
For a finite multilayer Maxwell body, $k_{2}(s)$ is a rational function of the Laplace variable and admits the mode decomposition (Gevorgyan et al., 2023, Eq. 26)
$k_{2}(s)=k_{E}+\sum_{i=1}^{n+1}\frac{r_{i}}{s-s_{i}},$ (2)
where $k_{E}$ is the elastic Love number, $s_{i}<0$ are the secular poles, $r_{i}$ are the corresponding residues, and $\tau_{i}=-s_{i}^{-1}$ are the relaxation times. Here $n$ denotes the number of Kelvin–Voigt elements of the equivalent homogeneous model, so the layered body has $n+1$ normal modes, in agreement with Gevorgyan et al. (2023, Eq. 22). Thus, the layered tidal response is fully characterized by the set of parameters $\{k_{E},s_{i},r_{i}\}$.
@@ 2.2 The homogeneous model
The equivalent homogeneous representation of stratified bodies is given by a generalized Voigt rheology in Figure 1. The Love number of the body is (Gevorgyan et al., 2023, Eq. 2)
$k_{2}(s)=\frac{3\mathrm{I}_{\circ}G}{R^{5}}\,C(s),$ (3)
where $\mathrm{I}_{\circ}$ is the mean moment of inertia, $C(s)$ is the complex compliance, and $R$ is the mean radius of the body. The gravitational rigidity $\gamma$ is defined from the fluid Love number $k_{f}$ as (Gevorgyan et al., 2023, Eq. 4)
$\gamma=\frac{3\mathrm{I}_{\circ}G}{R^{5}}\frac{1}{k_{f}}.$ (4)
For the generalized Voigt rheology, the complex compliance is decomposed as (Gevorgyan et al., 2023, Eq. 14)
@@
where $\alpha$ is the elastic rigidity and $C_{v}(s)$ contains the viscous contribution of the dashpot and a finite number of Kelvin–Voigt elements in Figure 1. The equivalent homogeneous model is specified by the parameters $\{\gamma,\alpha,\eta,\alpha_{i},\eta_{i}\}$. In these expressions $\gamma$, $\alpha$, and $\alpha_{i}$ have dimensions of $\mathrm{s^{-2}}$, as does the mapping constant $\Theta=3\mathrm{I}_{\circ}G/R^{5}$, while $\eta$ and $\eta_{i}$ have dimensions of $\mathrm{s^{-1}}$. To convert to SI units of Pa and $\mathrm{Pa\,s}$, we divide by the rescaling factor $\frac{152\pi}{15}\frac{R}{m_{1}}$ (Gevorgyan et al., 2020, Section 7.1.1), where $m_{1}$ is the mass of the body. The outputs reported by the package (for example in Table 6) are given in these units.
2.3 Extraction of the homogeneous model parameters
The parameters of the equivalent homogeneous generalized Voigt rheology are obtained from the layered pole–residue data $\{k_{E},s_{i},r_{i}\}_{i=1}^{n+1}$ by a purely algebraic procedure. Writing the layered compliance
$C_{L}(s)=\frac{R^{5}}{3\mathrm{I}_{\circ}G}\,k_{2}(s)=k_{E}^{C}+\sum_{i=1}^{n+1}\frac{r_{i}^{C}}{s-s_{i}},$ (6)
where $k_{E}^{C}$ and $r_{i}^{C}$ are $k_{E}$ and $r_{i}$ rescaled by $R^{5}/(3\mathrm{I}_{\circ}G)$, the elastic and gravitational parameters follow from the asymptotic limits of the compliance,
$\gamma=\left(k_{E}^{C}-\sum_{i=1}^{n+1}\frac{r_{i}^{C}}{s_{i}}\right)^{-1},\qquad\alpha=\frac{1}{k_{E}^{C}}-\gamma.$ (7)
The remaining parameters are $\eta$ and the $n$ Kelvin–Voigt pairs $\{(\alpha_{i},\eta_{i})\}_{i=1}^{n}$, totalling $2n+1$ unknowns. Attempting to extract these parameters from the $n+1$ pole conditions of the generalized Voigt model would be both underdetermined and nonlinear. Instead, we construct the complex compliance of the visco-elastic element $\hat{J}(s)$ (Gevorgyan et al., 2023, Eq. 13),
$\hat{J}(s)=\frac{1}{\alpha}+\frac{1}{\eta s}+\sum_{i=1}^{n}\frac{1}{\alpha_{i}+s\eta_{i}},$ (8)
directly from the layered data through the relation $C^{-1}(s)=\gamma+\hat{J}^{-1}(s)$, which yields
$\hat{J}(s)=\frac{C_{L}(s)}{1-\gamma\,C_{L}(s)}.$ (9)
All Kelvin–Voigt parameters are then read off from the partial-fraction decomposition of $\hat{J}(s)$.
Writing the layered compliance (6) as a single fraction $C_{L}(s)=N(s)/D(s)$, where $D(s)=\prod_{i=1}^{n+1}(s-s_{i})$ and $N(s)$ is a polynomial of degree $n+1$, equation (9) gives
$\hat{J}(s)=\frac{N(s)}{D(s)-\gamma\,N(s)}=\frac{N(s)}{s\,Q(s)},$ (10)
where the denominator vanishes at $s=0$ as a direct consequence of the definition of $\gamma$ in (7), and $Q(s)$ is a polynomial of degree $n$. The poles of $\hat{J}(s)$ are therefore $s=0$ and the $n$ roots of $Q(s)$, which are real, negative, and simple (from the interlacing property in Eq. 22 of Gevorgyan et al., 2023).
The residue of $\hat{J}(s)$ at $s=0$ comes entirely from the dashpot term in (8), the only singular contribution at the origin. Matching the residue computed from (10) gives
$\eta=\frac{Q(0)}{N(0)}.$ (11)
At each remaining pole $q_{i}$, only the $i$-th Kelvin–Voigt term is singular, and writing $1/(\alpha_{i}+s\eta_{i})=\eta_{i}^{-1}/(s-q_{i})$ with $q_{i}=-\alpha_{i}/\eta_{i}$, residue matching yields
$\eta_{i}=\frac{q_{i}\,Q^{\prime}(q_{i})}{N(q_{i})},\qquad\alpha_{i}=-q_{i}\eta_{i}=\frac{-q_{i}^{2}\,Q^{\prime}(q_{i})}{N(q_{i})}.$ (12)
No nonlinear system is solved. The only numerical step is the polynomial root-finding of $Q(s)$, a standard operation for a polynomial whose roots are real, negative, and simple. For a layered body with $n+1$ normal modes, the $2n+3$ data points $\{k_{E}^{C},r_{i}^{C},s_{i}\}$ uniquely determine the $2n+3$ parameters $\{\gamma,\alpha,\eta,\alpha_{i},\eta_{i}\}$ of the equivalent homogeneous model.
@@ 3.2 Input configuration
The single tidal forcing frequency $\omega$ and the frequency interval $\omega_{\min},\omega_{\max}]$ play distinct roles. The rheological parameters of the equivalent homogeneous model are frequency-independent, determined entirely by the secular poles and residues of the layered $k_{2}(s)$. The single $\omega$ is used only to evaluate and report $k_{2}(i\omega)$ at the frequency of interest for the body under consideration. For the Moon this corresponds to the orbital frequency $\omega=2.662\times 10^{-6}\penalty 10000\ \mathrm{rad\,s^{-1}}$, associated with the 27.32-day orbital period, which we use throughout the case study in Section 4. The interval $\omega_{\min},\omega_{\max}]$ enters only in the optional dominant-mode reduction step, where it defines the band over which each Kelvin–Voigt element is assessed for retention in the reduced model.
Table 1: Main user inputs and numerical controls of the Wolfram package. [^2] Physical quantities are supplied through ModelInput.wl; optional numerical settings control pole search and dominant-mode selection.
@@
rhok kg m-3 Layer densities
@@ 3.3 Main routine and computational workflow
performs the parameter extraction through the sequence of operations illustrated in Fig. 2. It follows the same three-step structure of Sections 1 and 2: forward computation of the layered Love number, mode decomposition, and rheological inversion. The code first evaluates the elastic Love number $k_{E}$ and the complex tidal response $k_{2}(i\omega)$ for the layered body. It then searches for the zeros of the secular determinant on the negative real axis, refines them numerically, and computes the corresponding residues. From these quantities it evaluates the fluid Love number $k_{f}$ and the mapping constant $\Theta=3\mathrm{I}_{\circ}G/R^{5}$, and applies the algebraic procedure of Section 2.3 to extract the parameters $\{\gamma,\alpha,\eta,\alpha_{i},\eta_{i}\}$ of the equivalent generalized Voigt model.
@@
The forward layered problem is evaluated through a unified $6\times 6$ propagator matrix approach for all layers. Liquid layers are treated by small-shear regularization, such that solid and liquid regions are handled within the same numerical framework. The current package does not switch to a separate Clairaut propagator for liquids, but instead uses a single propagation formalism for arbitrary combinations of solid and liquid layers. A true liquid has $\mu=0$, but the $Y$-matrix (Sabadini et al., 2016, Eq. 2.42) contains terms of the form $\mu/r$ in rows 3, 4, and 6. Setting $\mu=0$ exactly would therefore make these rows degenerate, causing the linear solver to fail when computing the propagator
@@
Instead, we assign a very small effective rigidity $\hat{\mu}\approx 10^{-20}\,\mathrm{Pa}$ to liquid layers. This keeps the matrix full-rank and invertible, while making the shear-stress rows effectively zero to machine precision. This value is negligible compared with physical shear moduli, which are typically of order $10^{9}$$10^{11}\,\mathrm{Pa}$, but remains large enough to preserve numerical stability at all reasonable (actually, even at quite stringent) working precisions. This is a standard regularization in viscoelastic normal-mode calculations that avoids introducing a separate $2\times 2$ or $4\times 4$ propagator for liquid layers while yielding results indistinguishable from the inviscid limit to within working precision.
After the full generalized Voigt model is obtained, the package performs an optional dominant-mode reduction, following the criterion introduced in Gevorgyan et al. (2023, Appendix A). For each Voigt element, it evaluates the maximum fractional contribution to $\operatorname{Im}[C(i\omega)]$ over a user-defined frequency interval $\omega_{\min},\omega_{\max}]$. Elements whose contribution exceeds the prescribed threshold are kept in the reduced model. This produces a compact, effective rheology targeted to the frequency interval most relevant for the intended dynamical application.
@@ 3.4 Output
The output falls into three groups: the Love numbers of the layered body, the constants used in the layered-to-homogeneous mapping, and the relaxation modes and rheological parameters of the equivalent generalized Voigt model. This organization separates diagnostic quantities from the final effective model. The Love-number and mode outputs are used to check the inversion, and the generalized Voigt parameters are used in downstream applications. The package also generates summary tables, dominant-mode diagnostics, and comparison plots. The user can inspect the mode decomposition, compare the full and reduced models, or export the reduced generalized Voigt parameters to an external integrator such as RheoVolution (de Oliveira et al., 2025a, b).
Table 2: Quantities returned by computeHomogeneous. [^3] The routine returns both the intermediate mode information of the layered model and the parameters of the equivalent homogeneous generalized Voigt rheology.
@@
prefactor s-2 Mapping prefactor $\Theta=3\mathrm{I}_{\circ}G/R^{5}$
@@
nPoles integer Number of secular poles (equal to $n+1$ in the notation of Section 2.3)
@@
nVoigt integer Number of Kelvin–Voigt elements (equal to $n$ in the notation of Section 2.3)
@@ 4 Case study: A five-layer model of the Moon
To test the package, we use the five-layer lunar model of Gevorgyan et al. (2023) to illustrate the equivalence between a multilayer Maxwell body and an effective homogeneous rheology. The model parameters are summarized in Table 3, and its cross-section is illustrated in Fig. 3; the figure was automatically produced by the VisualizeLayers.wl module. The Moon is the default example configuration in the package. The package was also tested for Io, Europa, and Enceladus using models and data from Hussmann and Spohn (2004), Wahr et al. (2006), Bierson and Nimmo (2016), and Thomas et al. (2016), although we are only going to report the results for the Moon. The ModelInput files for these additional test cases can be found in the software package repository.
Table 3: Physical parameters for the 5-layer viscoelastic model of the lunar interior. [^4] Values taken from Matsuyama et al. (2016, Table 2) and Gevorgyan et al. (2023).